# A Unified Model for Multi-epoch Neutrino Events and Broadband Spectral Energy Distribution of TXS 0506+056

### 12 Jan 2024 v1

@@ 1 Introduction
The origin of high-energy neutrinos is still unclear (Aartsen et al., 2013; IceCube Collaboration, 2013). As the most powerful persistent emitters of electromagnetic (EM) radiation in the universe, blazars have long been suggested as potential accelerators of high-energy cosmic-ray protons (or nuclei) and sources of high-energy neutrinos (e.g., Mannheim, 1993, 1995; Stecker & Salamon, 1996; Halzen & Hooper, 2002; Atoyan & Dermer, 2001; Murase et al., 2014). Such speculations are supported by the discovery of possible correlation between high-energy neutrino events and blazars (IceCube Collaboration et al., 2018a, b). These events have been extensively studied in the framework of the so-called photohadronic interaction model (Ansoldi et al., 2018; Keivani et al., 2018; Murase et al., 2018; Padovani et al., 2018, 2019; Cerruti et al., 2019; Gao et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Xue et al., 2019; Abbasi et al., 2021) and hadronuclear interaction model (Sahakyan, 2018; Liu et al., 2019; Banik & Bhadra, 2019). In these hadronic processes, the produced secondary EM particles can induce EM cascades, contributing to EM radiation from the optical band to the X-ray band or even to the MeV gamma-ray band (Xue et al., 2021; Reimer et al., 2019). On the other hand, the low-energy radiation of a blazar can be the target radiation field for the photohadronic interactions (Kelner & Aharonian, 2008; Murase et al., 2014; Murase & Stecker, 2022). In this sense, the neutrino emission from a blazar is related to its broadband radiation.
@@ 3 MODEL SETUP AND METHOD
To research the neutrino events of TXS 0506+056 in different epochs, we model the broadband SED in each epoch, based on which we can calculate corresponding neutrino detection rate in the epoch. Similar to the standard one-zone model for the blazar jet, we consider the electron synchrotron radiation, the inverse Compton scattering off the synchrotron radiation field (i.e., synchrotron self-Compton, SSC, Harris & Krawczynski 2006) and off the external radiation (i.e., external Compton, EC) of electrons. We also take into account the proton synchrotron radiation, the Bethe-Heitler process, and the photopion process, as well as the emission of pairs generated in the electromagnetic cascade initiated by these processes.
It is assumed that relativistic particles are injected into a blob (flaring zone) which has a spherical geometry with a radius $R$, filled with a uniformly entangled magnetic field $B$. We assumed the jet’s bulk Lorentz factor $\Gamma$ remains constant up to 0.1 pc and we approximate the Doppler factor by $\delta=\Gamma$, $\delta_{0}$ is the Doppler factor at 0.1 pc. Considering a truncated conical jet structure and a dissipation zone of radius $R$ comparable to the transverse radius of the jet at certain distance $r$ from the SMBH (Wang et al., 2022), we have
@@
Relativistic electrons or protons are usually assumed to be injected in the blob with a broken power-law distribution (Ghisellini et al., 2010; Wang et al., 2020) or a power-law distribution, i.e.,
@@
where $t_{\rm e(p)}=$min$\left\{t_{\rm cool},t_{\rm e(p),dyn}\right\}$. For electrons $t_{\rm cool}=3m_{\rm e}c/\left({4\left(U_{\rm B}+\kappa_{\rm KN}U_{\rm ph}\right)\sigma_{\rm T}\gamma_{\rm e}}\right)$ is the electron radiative cooling timescale, with $U_{\rm B}$ being the energy density of the magnetic field, and $U_{\rm B}=B^{2}/8\pi$. $U_{\rm ph}$ being the energy density of the soft photons. $\sigma_{T}$ is the Thomson scattering cross section and $\kappa_{\rm KN}$ is a numerical factor accounting for the Klein-Nishina effect(Moderski et al., 2005). For protons, we need to include the energy loss caused by the photopion production and the Bethe-Heitler (BH) pair production when calculating the cooling timescale. These two processes will be further described later. $t_{\rm e(p),dyn}=R/c$ is the the dynamical timescale of electrons or protons, here $R=r\sin\theta$ is the radius of the blob, $r$ is the distance between the black hole and the blob. All timescales are evaluated in the comoving frame of the blob.
Under the unified model, the synchrotron, SSC and EC radiations are calculated following Wang et al. (2022). If the blob (flaring zone) is close to the supermassive black hole (SMBH), the radiation of the broad line region (BLR) and the dust torus (DT) could enter the blob with a significant amplification of the energy density due to the Doppler effect. The BLR and DT radiation is taken as an isotropic greybody with a peak at $\approx 2\times 10^{15}\Gamma\rm\ Hz$ (Tavecchio & Ghisellini, 2008) and $3\times 10^{13}\Gamma\rm\ Hz$ (Cleary et al., 2007) in the jet comoving frame, respectively. The energy density of BLR ($u_{\rm BLR}$) and DT ($u_{\rm DT}$) emission can be approximated (Hayashida et al., 2012) by
@@
where $r$ is the distance between the central black hole and the dissipation region. $\eta_{\rm BLR}=0.1$ and $\eta_{\rm DT}=0.1$ are the fractions of the disk luminosity $L_{\rm d}$ reprocessed into radiations of BLR and DT, respectively, $r_{\rm BLR}=0.1\left(L_{\rm d}/10^{46}\rm erg\ s^{-1}\right)^{1/2}\rm pc$ and $r_{\rm DT}=2.5\left(L_{\rm d}/10^{46}\rm erg\ s^{-1}\right)^{1/2}\,$pc are the characteristic distances where the above processes taking place. The disk luminosity $L_{\rm d}$ is assumed to be $5\times 10^{44}\rm erg\ s^{-1}$ (Xue et al., 2021). Meanwhile, the X-ray photons emitted by the corona surrounding the accretion disk (Heckman & Best, 2014) are also considered in our model, if the blob is near or at the jet base with a distance comparable to only a few times larger than the Schwarzschild radius ($r_{\rm Sch}\sim 10^{14}(M_{\rm SMBH}/3\times 10^{8}M_{\odot})\rm cm$) of the central SMBH, the X-ray photons would interact with gamma-ray photons emitted by the blob (Righi et al., 2019). The emission of the corona typically has a power-law spectrum (Ricci et al., 2018; Xue et al., 2021)
@@
here we simply assume that its spectral index $\alpha=1$ and $L_{1\ \rm keV}$ is set as $5\times 10^{43}\rm erg\ s^{-1}$ for TXS 0506+056 (Xue et al., 2021). The corresponding energy density in the comoving frame can be estimated as
@@
Note that these radiation field can also interact with relativistic protons via the Bethe–Heitler pair-production and photopion production. The photopion process is the production of pions in proton-photon interactions:
@@
via which unstable pions are produced and further decay into
@@
Another important process is the Bethe–Heitler pair-production, leading to the production of electron/positron pairs, i.e.,
@@
We follow the semi-analytical method developed by Kelner & Aharonian (2008) calculate spectra of secondary particles in these two processes. The cooling timescales of both electrons and protons via the aforementioned processes in the model, which will be discussed in the following section, are shown in Figure 1 for references.
The maximum protons Lorentz factor in the emission region can be approximated by equating the acceleration and the cooling or dynamical timescales
@@
where $\alpha(\geq 1)$ is the parameter characterising the particle acceleration efficiency ($\alpha=1$ being the most efficient case), which depends on the spectrum of magnetic turbulence and on the velocity of the scattering center in the case of Fermi-type acceleration (Sikora, 2011). Here we employ $\alpha=10$ in our model. The kinetic luminosity of the magnetic field and nonthermal particles can be approximated as (Celotti & Ghisellini, 2008)
@@
where $U_{i}$ is the energy density of magnetic field or nonthermal particles, here i = $\left\{\rm B,\rm e,\rm p\right\}$, represents magnetic field, electrons and protons respectively.
Figure 1: The cooling timescale of protons and electrons in different epochs, as well as the opacity of the intrinsic $\gamma\gamma$ absorption of the radiation zone. The horizontal axis $\rm eV$ represents ernergy of protons for synchrotron, photopion and Bethe-Heitler process in the comoving frames, and also represents ernergy of electrons for synchrotron, SSC and EC process in the comoving frames. In addition, we also plot the optical depth of photons(the red dot-dashed line) in the figure. The vertical grey dashed lines show the maximum electron and proton energy allowed by the $t_{\rm acc}=$min$\left\{t_{\rm cool},t_{\rm e(p),dyn}\right\}$. Parameters of each model are described in Table 1, respectively.
@@ 4 SED Modeling and results
In this section, we use the stochastic dissipation model to reproduce the mutiwavelength emission of TXS 0506+056 in the four epochs with neutrino detection. For the photopion process and the BH process, radiation fields of the blazar serve as targets, including the emission of the BLR, the dusty torus and the corona, the radiation of primary electrons, as well as the radiation of secondary pairs developed in the electromagnetic cascade initiated by the two processes. The photopion process and the BH process including the cascade are calculated following the method shown in (Kelner & Aharonian, 2008; Böttcher et al., 2013).

### 18 Jan 2024 v3

@@ 1 Introduction
The origin of high-energy neutrinos is still unclear (Aartsen et al., 2013; IceCube Collaboration, 2013). As the most powerful persistent emitters of electromagnetic (EM) radiation in the universe, blazars have long been suggested as potential accelerators of high-energy cosmic-ray protons (or nuclei) and sources of high-energy neutrinos (e.g., Mannheim, 1993, 1995; Stecker & Salamon, 1996; Halzen & Hooper, 2002; Atoyan & Dermer, 2001; Murase et al., 2014). The discovery of a possible correlation between high-energy neutrino events and blazars supports this speculation. (IceCube Collaboration et al., 2018a, b). Many researchers have studied these events in the framework of the so-called photohadronic interaction model (Ansoldi et al., 2018; Keivani et al., 2018; Murase et al., 2018; Padovani et al., 2018, 2019; Cerruti et al., 2019; Gao et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Xue et al., 2019; Abbasi et al., 2021) and hadronuclear interaction model (Sahakyan, 2018; Liu et al., 2019; Banik & Bhadra, 2019). In these hadronic processes, the produced secondary EM particles can induce EM cascades, contributing to EM radiation from the optical band to the X-ray band or even to the MeV gamma-ray band (Xue et al., 2021; Reimer et al., 2019). On the other hand, the low-energy radiation of a blazar can be the target radiation field for the photohadronic interactions (Kelner & Aharonian, 2008; Murase et al., 2014; Murase & Stecker, 2022). In this sense, the neutrino emission from a blazar is related to its broadband radiation.
@@ 3 MODEL SETUP AND METHOD
To research the neutrino events of TXS 0506+056 in different epochs, we model the broadband SED in each epoch, based on which we can calculate corresponding neutrino detection rate in the epoch. In our model we consider the electron synchrotron radiation, the inverse Compton scattering off the synchrotron radiation field (i.e., synchrotron self-Compton, SSC, Harris & Krawczynski 2006) and off the external radiation (i.e., external Compton, EC) of electrons. We also take into account the proton synchrotron radiation, the photopion process, and the Bethe-Heitler process, as well as the emission of pairs generated in the electromagnetic cascade initiated by these processes.
We assumed that the flaring zone is a blob which has a spherical geometry with a radius $R$, filled with a uniformly entangled magnetic field $B$. The jet’s bulk Lorentz factor $\Gamma$ is assumed to remain constant up to 0.1 pc and we approximate the Doppler factor by $\delta=\Gamma$, $\delta_{0}$ is the Doppler factor at 0.1 pc. Considering a truncated conical jet structure and a dissipation zone of radius $R$ comparable to the transverse radius of the jet at certain distance $r$ from the SMBH (Wang et al., 2022), we have
@@
We assumed that relativistic electrons or protons are injected in the blob with a broken power-law distribution (Ghisellini et al., 2010; Wang et al., 2020) or a power-law distribution, i.e.,
@@
where $t_{\rm e(p)}$ is the minimum value among $t_{\rm cool}$ and $t_{\rm e(p),dyn}$. $t_{\rm cool}=3m_{\rm e}c/\left({4\left(U_{\rm B}+\kappa_{\rm KN}U_{\rm ph}\right)\sigma_{\rm T}\gamma_{\rm e}}\right)$ is the radiative cooling timescale of electron. $U_{\rm B}=B^{2}/8\pi$ is the energy density of the magnetic field and $U_{\rm ph}$ is the energy density of the soft photons. Here $\sigma_{T}$ is the Thomson scattering cross section and $\kappa_{\rm KN}$ is a factor accounting for the Klein-Nishina effect(Moderski et al., 2005). For protons, we need to include the energy loss caused by the photopion production and the Bethe-Heitler (BH) pair production when calculating the cooling timescale. These two processes will be further described later. $t_{\rm e(p),dyn}=R/c$ is the electron(proton) dynamical timescale, here we assumed the radius of the blob is $R=r\sin\theta$, $r$ is the distance between the blob and the black hole. In this work all timescales are evaluated in the comoving frame of the blob.
Under the unified model, the synchrotron, SSC and EC radiations are calculated following Wang et al. (2022). If the blob (flaring zone) is close to the supermassive black hole (SMBH), the radiation of the broad line region (BLR) and the dust torus (DT) could enter the blob with a significant amplification of the energy density due to the Doppler effect. The radiation field of BLR is assumed to be a blackbody emission peaking at a frequency $2\times 10^{15}\Gamma\rm\ Hz$ (Tavecchio & Ghisellini, 2008), and the radiation field of DT is assumed to be a blackbody emission peaking at a frequency $3\times 10^{13}\Gamma\rm\ Hz$ (Cleary et al., 2007) in the jet comoving frame, respectively. The energy density of external photons from the BLR and DT can be approximated (Hayashida et al., 2012) by
@@
where $\eta_{\rm BLR}=\eta_{\rm DT}=0.1$ are the fractions of the disk luminosity $L_{\rm d}$ reprocessed into the BLR and DT radiation, respectively. We assumed the characteristic distance of BLR in the AGN frame is $r_{\rm BLR}=0.1\left(L_{\rm d}/10^{46}\rm erg\ s^{-1}\right)^{1/2}\rm pc$ and the characteristic distance of DT is $r_{\rm DT}=2.5\left(L_{\rm d}/10^{46}\rm erg\ s^{-1}\right)^{1/2}\,$pc. The disk luminosity $L_{\rm d}$ is assumed to be $5\times 10^{44}\rm erg\ s^{-1}$ (Xue et al., 2021). Meanwhile, the X-ray photons emitted by the corona surrounding the accretion disk (Heckman & Best, 2014) are also considered in our model, if the blob is near or at the jet base with a distance comparable to only a few times larger than the Schwarzschild radius ($r_{\rm Sch}\sim 10^{14}(M_{\rm SMBH}/3\times 10^{8}M_{\odot})\rm cm$) of the central SMBH, the X-ray photons would interact with gamma-ray photons emitted by the blob (Righi et al., 2019). We assumed the emission of the corona has a power-law spectrum (Ricci et al., 2018; Xue et al., 2021)
@@
here we simply assume that its spectral index $\alpha=1$ and $L_{1\ \rm keV}$ is set as $5\times 10^{43}\rm erg\ s^{-1}$ for TXS 0506+056 (Xue et al., 2021). The corresponding energy density in the comoving frame can be written as
@@
Note that the relativistic protons can also interact with these radiation field via the photopion process and Bethe–Heitler process. The photopion process is the production of pions in proton-photon interactions:
@@
via which unstable pions are produced and pions will decay into
@@
Another important process is the Bethe–Heitler process, this process will lead to the production of electron/positron pairs, i.e.,
@@
Kelner & Aharonian (2008) developed a semi-analytical method to calculate spectra of secondary particles in these two processes and we followed the method in Kelner & Aharonian (2008). The cooling timescales of both electrons and protons via the aforementioned processes in the model are shown in Figure 1 for references.
The maximum Lorentz factor of protons in the emission region can be approximated by equating the acceleration and the cooling or dynamical timescales
@@
where $\alpha(\geq 1)$ is the parameter characterising the particle acceleration efficiency ($\alpha=1$ being the most efficient case), which depends on the spectrum of magnetic turbulence and on the velocity of the scattering center in the case of Fermi-type acceleration (Sikora, 2011). Here we employ $\alpha=10$ in our model. The kinetic luminosity of the nonthermal particles and magnetic field can be written as (Celotti & Ghisellini, 2008)
@@
where $U_{i}$ is the energy density of nonthermal particles or magnetic field, here i = $\left\{\rm e,\rm p,\rm B\right\}$, represents electrons, protons and magnetic field respectively.
Figure 1: The cooling timescale of protons and electrons in different epochs, as well as the opacity of the intrinsic $\gamma\gamma$ absorption of the radiation zone. The horizontal axis $\rm eV$ represents ernergy of protons for synchrotron, photopion and Bethe-Heitler process in the comoving frames, and also represents ernergy of electrons for synchrotron, SSC and EC process in the comoving frames. In addition, we also plot the optical depth of photons(the red dot-dashed line) in the figure. The vertical grey dashed lines show the maximum electron and proton energy allowed by the $t_{\rm acc}=$min$\left\{t_{\rm cool},t_{\rm e(p),dyn}\right\}$. Parameters of the model are described in Table 3, respectively.
@@ 4 SED Modeling and results
In this section, we use the stochastic dissipation model to reproduce the mutiwavelength emission of TXS 0506+056 in the four epochs with neutrino detection. For the photopion process and the BH process, radiation fields of the blazar serve as targets, including the emission of the BLR, the dusty torus and the corona, the radiation of primary electrons, as well as the radiation of secondary pairs developed in the electromagnetic cascade initiated by the two processes. The BH process and photopion process including the cascade initiated by these processes are calculated following the method shown in (Kelner & Aharonian, 2008; Böttcher et al., 2013).